# Probing the accretion geometry of black hole X-ray binaries: A multi-mission spectro-polarimetric and timing study

### v1

@@ 4.4 Energy-dependent Polarization
Figure 5: Variation of polarization degree (PD) with energy. In each panel, source name and spectral state (observed during different epoch) are marked. The histograms in colors represent the minimum detectable polarization at $99\%$ confidence ($\rm MDP_{99}\%$). See the text for details.
@@
We further explore energy-dependent polarization properties of the selected sources that confirm the presence of significant polarized emission in $2-8$ keV energy range of IXPE. For each source except 4U $1957+115$, IGR J17091362417091362417091-362417091 - 3624 and GX $339-4$, we divide $2-8$ keV energy range into six linearly spaced energy bins and estimate polarization using the model-independent approach (see also Svoboda et al. 2024a; Veledina et al. 2023). For 4U $1957+115$, IGR J17091362417091362417091-362417091 - 3624 and GX $339-4$, only four energy bins are considered for the energy-dependent study to ensure relatively improved photon statistics per energy bin (see also Marra et al. 2024). Further, we estimate the significance of the energy variation of PD over these bins following Krawczynski et al. (2022); Majumder et al. (2024b). In doing so, we consider PD in $2-8$ keV energy range (see Table 5) as the null hypothesis against which the significance of energy variations in PD is computed for each epoch of a given source. With this, following Majumder et al. (2024b), we calculate the probability of rejecting the null hypothesis using $(N-1)$ degrees of freedom, taking into account the variation of PD over $N$ energy bins for the respective cases. This essentially indicates the confidence level at which the PD variation is significant over different energy bins. The obtained significance levels of the energy variation of PD are mentioned in Table 5 for the respective cases. We observe that this significance varies from $\lesssim 1\sigma$ to $>4\sigma$ across different epochs of the sources in various spectral states. In Fig. 5, we present the variation of PD with energy over different observational epochs using filled circles in the respective panels. The histograms in colors represent the $\rm MDP_{99}\%$ associated with the measurements in the selected energy bins for the corresponding epochs. Note that, for a given source, when multi-epoch observations are available, we present the results for two epochs of different spectral states only, except LMC X--3.
@@
Interestingly, we observe that Cyg X11-1- 1 and Swift J1727.816131613-1613- 1613 show similar energy variation of PD within $\sim 2-6\%$ in different spectral states (LHS, HSS, HIMS and SIMS). Moreover, for Cyg X11-1- 1, the variation in PD with energy remains most significant at a confidence level exceeding $4\sigma$ in HSS, whereas it remains $2.8\sigma$ for Swift J1727.816131727.816131727.8-16131727.8 - 1613 in HIMS, when all observational epochs are considered (see Fig. 5). Note that, PD $<\rm MDP_{99}\%$ at higher energy bins in a few epochs of several sources, indicates a null-detection. Interestingly, unlike Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613, a weak energy dependency of PD is noticed for IGR $17091-3624$ in the LHS, for which PD corresponding to two out of four energy bins remain below the $\rm MDP_{99}\%$ (see Fig. 5). We find a significant increase in PD with energy up to $\sim 11\%$ in HSS ($>4\sigma$) and SPL ($>4\sigma$) state of 4U 16304747-47- 47 (see Fig. 5). Similarly, distinct variations of PD over several energy bins are observed for Cyg X33-3- 3 during LHS ($>4\sigma$) and SIMS ($>4\sigma$). LMC X33-3- 3 shows constant polarization over the initial three energy bins with a sudden jump in PD up to $\sim 7.8\%$ at $\sim 5.4$ keV in both epochs (see Fig. 5). Further, we only observe a marginal variation ($1.7\sigma$) of PD in 4U 1957$+115$ up to $\sim 3\%$. Interestingly, significant ($3.3\sigma$) energy variation of polarization degree is observed for GX $339-4$. Although, PD $<\rm MDP_{99}\%$ is obtained in all four energy bins (see Fig. 5). Overall, we state that a significant ($>3\sigma$) variation of polarization degree with energy is observed only for Cyg X11-1- 1, 4U 16304747-47- 47 and Cyg X33-3- 3, whereas a moderate ($2\sigma$ to $3\sigma$) variation is seen for Swift J1727.816131613-1613- 1613 and LMC X33-3- 3. However, Veledina et al. 2024a, b reported an energy-independent variation of PD across 12 energy bins for Cyg X--3. We mention that the strong ($>4\sigma$) energy dependence of PD observed in our analysis possibly resulted due to relatively small uncertainties obtained in PD estimates over the energy bins and a significant drop in PD around the unpolarized iron line complex at $\sim 6.4$ keV. It is worth mentioning that the energy variation of PA for all sources under consideration remains statistically insignificant ($<2\sigma$). However, we note that PA drops from $-30^{\circ}$ to $-60^{\circ}$ at $\sim 6$ keV in 4U $1957+115$, which is marginally significant at $1.6\sigma$ confidence level.
@@ 4.5 Spectro-polarimetric Modeling
We explore the model-dependent spectro-polarimetric properties of the sources from the simultaneous fitting of the I, Q and U Stokes spectra of IXPE in $2-8$ keV energy range. Due to the calibration issues Krawczynski et al. (2022); Veledina et al. (2024a); Steiner et al. (2024), only DU1 data is considered for the spectro-polarimetric study of Cyg X11-1- 1, Cyg X33-3- 3 and Swift J1727.816131613-1613- 1613, whereas data combining all the DUs are used for the remaining sources. For simplicity, we first adopt the model combination M4: Tbabs×$\times$×polconst×$\times$(diskbb + powerlaw) for the spectro-polarimetric fitting. Here, polconst represents a constant polarization model, where the degree and angle of polarization are treated as the model parameters. The best-fit results yield $\chi_{\rm red}^{2}$ ranging from $0.91-1.79$ across all selected sources, with PD varying between $\sim 1.16-13.06\%$ and PA spanning from $-89.13^{\circ}$ to $85.15^{\circ}$. Note that the spectro-polarimetric modeling of Swift J1727.816131727.816131727.8-16131727.8 - 1613 does not constrain the hydrogen column density, likely due to low-energy threshold of IXPE ($\sim 2$ keV; see Veledina et al., 2023; Ingram et al., 2024).
Next, we investigate the energy-dependent polarization properties using model-dependent approach for all the sources that demonstrate the signature of X-ray polarization. In doing so, we replace the polconst component with the energy-dependent polarization model polpow [^15] https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node218.html. Empirically, polpow assumes a power-law variation in the polarization parameters of the form PD($E$) $=A_{\rm norm}\times E^{-A_{\rm index}}$ (in fraction) and PA($E$) $=\psi_{\rm norm}\times E^{-\psi_{\rm index}}$ (in ) while modeling the Q and U Stokes spectra. We find that the best-fitted $\psi_{\rm index}$ remains consistent with zero within its $1\sigma$ errors and hence, we freeze it to zero during the spectral fitting of all observations. Accordingly, we obtain the best spectral fit for all sources with $\chi_{\rm red}^{2}\sim 0.97-1.46$. Furthermore, integrating PD($E$) in $2-8$ keV energy range using the best-fitted model parameters ($A_{\rm index}$ and $A_{\rm norm}$), we obtain PD as $3.23-14.33\%$. Moreover, with $\psi_{\rm index}$ frozen to zero, the polarization angles are obtained as the best-fitted $\psi_{\rm norm}$ values that vary from $-88.92^{\circ}$ to $85.11^{\circ}$. Note that, in a few observations of Cyg X11-1- 1, 4U 1957$+115$ and GX $339-4$ (see Table 6), the data is unable to constrain the parameters of polpow component even at $1\sigma$ level, and hence, we refrain from the modeling with polpow for these observations. In Table 6, we present the best-fitted model parameters obtained from the spectro-polarimetric study of all sources. We find that the model-dependent polarimetric results are in agreement with the findings of the model-independent study except Cyg X33-3- 3, confirming the robustness of the polarization measurements for the individual sources. However, for Cyg X--3, we obtain significantly low PD compared to the results of the model-independent study, likely due to multiple low-energy unpolarized emission lines including strong iron complex around $\sim 6.4$ keV (see Veledina et al. 2024a), dominating the Stokes spectra.
@@ 4.6 Spectro-polarimetric Correlation
Figure 6: Variation of PD ($2-8$ keV), obtained from model-independent analysis, as a function of the disc-to-Comptonized flux ratio ($F_{\rm ratio}$) derived from wide-band spectral modelling. The color bar represents the covering fraction ($cov_{\rm frac}$), while each data point is labeled with the corresponding spectral state and observational epoch. Data points within the shaded region (light blue) are used to compute the Pearson correlation coefficient. Downward arrows indicate null-detections of polarization for the respective sources. See the text for details.
We examine the spectro-polarimetric correlation properties for each source based on the results obtained from wide-band spectral modeling using model M1 (see §4.2) and the polarimetric measurements (see §4.3). Specifically, we explore the correlation between the observed PD obtained from model-independent study, and best-fitted spectral parameters, namely the ratio of disc to Comptonized spectral flux ($F_{\rm ratio}$) in $1-100$ keV energy range, and the covering fraction ($cov_{\rm frac}$). In Fig. 6, we present the variation of PD with $F_{\rm ratio}$ (see Table 3) using different markers, where the covering fraction ($cov_{\rm frac}$, in percent) is represented by the colors of respective markers. The color bar on the right indicates the range of $cov_{\rm frac}$. The spectral state and the corresponding epoch (in parentheses) for each observation are marked in the figure.
@@
Furthermore, we find a strong correlation (anti-correlation) between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$) with $\rho\sim 0.9$ ($-0.9$) for Swift J1727.816131727.816131727.8-16131727.8 - 1613. In particular, a sharp drop in PD is noticed from $\sim 4.7\%$ (HIMS) to $\sim 0.4\%$ (HSS) (see also Svoboda et al. 2024b) during which $F_{\rm ratio}$ increases from $\sim 0.7$ to $\sim 66$ and $cov_{\rm frac}$ noticeably reduces down to $13\%$ from a maximum value of $82\%$. A similar trend is observed for Cyg X11-1- 1 with $\rho=0.9$ ($-0.9$) for the respective positive (negative) correlations between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$). As before, a low (high) $F_{\rm ratio}\sim 0.3$ ($11.5$) and high (low) $cov_{\rm frac}\sim 61-77\%$ ($0.5-31\%$) are observed in the LHS (HSS) of Cyg X11-1- 1 with PD $\sim 3.7-4.8\%$ ($1.4-2.1\%$) (see Fig. 6). Moreover, we also find a similar behavior in the LHS of IGR J17091362417091362417091-362417091 - 3624, for which a very high PD ($\sim 9\%$) is seen with $cov_{\rm frac}\sim 87\%$ and $F_{\rm ratio}\sim 0.2$. These findings suggest that a significant contribution in the polarization degree comes from the Comptonized emission for Swift J1727.816131727.816131727.8-16131727.8 - 1613, Cyg X11-1- 1 and IGR J17091362417091362417091-362417091 - 3624 in the HIMS and/or LHS. Furthermore, we notice a low polarization degree (PD $\lesssim 2\%$) for all the sources except 4U 16304747-47- 47 and LMC X33-3- 3 in the HSS/SIMS with $F_{\rm ratio}\gtrsim 2$ and $cov_{\rm frac}\lesssim 30\%$. Interestingly, 4U 16304747-47- 47 being the source of extreme soft nature ($F_{\rm ratio}\sim 52$, $cov_{\rm frac}\sim 5\%$) manifests a high polarization degree of $\sim 8.3\%$ in the HSS, whereas it reduces to $6.8\%$ in the SPL state with $F_{\rm ratio}\sim 3.6$ and $cov_{\rm frac}\sim 38\%$. We observe that 4U $1957+115$ (GX $339-4$) shows PD $\sim 1.2\%$ ($1.9\%$) with $cov_{\rm frac}\sim 2.7\%$ ($34\%$) and $F_{\rm ratio}\sim 18$ ($2.4$) in the HSS (SIMS). Moreover, LMC X33-3- 3 exhibits moderate polarization with PD in the range of $2.4-3\%$ for $F_{\rm ratio}\gtrsim 4$ and $cov_{\rm frac}\sim 0.5\%$ during the HSS.
@@ 5 Discussion
In this study, we present a comprehensive spectro-polarimetric investigation of eleven BH-XRBs using quasi-simultaneous observations from IXPE, NICER, AstroSat and NuSTAR. IXPE data are analyzed to examine the polarization properties in $2-8$ keV energy range, while combined NICER, AstroSat and NuSTAR observations provide insights into the spectro-temporal behavior across the wide energy range ($0.5-100$ keV). Subsequently, the spectro-polarimetric findings are used to investigate the accretion-ejection dynamics and geometry surrounding the sources.
@@ 5.1 Spectro-temporal Properties and Spectral States
Spectro-temporal studies using NICER, NuSTAR and AstroSat observations in $0.5-100$ keV energy range impart distinct variability properties in various spectral states of the sources (see §4.1 and §4.2). The obtained results are in line with the spectro-temporal characteristics generally seen in the respective spectral states of BH-XRBs (Belloni et al., 2005; Remillard & McClintock, 2006; Kalemci et al., 2006; McClintock et al., 2009; Nandi et al., 2012; Radhika & Nandi, 2014; Radhika et al., 2016; Sreehari et al., 2019b; Connors et al., 2019; Méndez et al., 2022; Athulya et al., 2022; Aneesha et al., 2024; Shui et al., 2024; Li et al., 2025). These findings reinforce the presence of distinct spectral states as observed during the multi-mission campaigns of the sources.
@@
In particular, we observe a high fractional variability amplitude $rms_{\rm tot}\%\sim 30-52$ in the PDS of LHS, whereas significantly low variability with $rms_{\rm tot}\%\sim 6-14$ are seen in the intermediate states (HIMS and SIMS) of the sources (see Table 3). Negligible power distribution at lower frequencies results in marginal variability ($rms_{\rm tot}\%\sim 1-10$) for the sources in the HSS (see Table 3 and Fig. 2). In addition, we find the presence of a strong type-C QPO feature at $\sim 1.4$ Hz (E1), which evolves to $\sim 6.7$ Hz (E2-E5), during the HIMS and SIMS of Swift J1727.816131613-1613- 1613 over a wide energy range of $0.5-100$ keV with NICER, NuSTAR and AstroSat observations (see Fig. 2). The origin of this QPO is attributed to a strong Comptonizing corona surrounding the central source (Chakrabarti & Titarchuk, 1995; Chakrabarti et al., 2008; Nandi et al., 2024); however, based on the correlation between the evolution of QPO frequency and quasi-simultaneous radio jet ejections, Liao et al. (2024) suggest a jet-like, vertically elongated corona for Swift J1727.816131727.816131727.8-16131727.8 - 1613. A similar type-C QPO at $\sim 0.2$ Hz is detected in LHS of IGR J17091362417091362417091-362417091 - 3624 (see Fig. 2). Moreover, we also observe a type-B QPO in the disc-dominated SIMS of GX $339-4$ in the $0.5-80$ keV energy range from NICER and AstroSat data. Similar type-B QPOs have been observed during $2024$ and earlier outbursts of GX $339-4$, indicating the presence of a compact corona located very close to the black hole, accompanied by dominant disc emission (Nandi et al., 2012; Aneesha et al., 2024).
@@
Admittedly, the wide-band energy spectral analysis ($0.7-70$ keV) reveals distinct spectral states of each source. The observed spectra of all sources are well described by a standard accretion disc component (diskbb) combined with thermal Comptonization of soft photons by a spherical corona (thcomp) (see Table 3 and Fig. 2). The spectral properties exhibit a strong correlation with the spectral states of the sources. In particular, during LHS and HIMS, the spectra are dominated by Comptonized emission, contributing $\sim 64-87\%$ of the total flux, accompanied by a high covering fraction ($cov_{\rm frac}\sim 61-87$). In contrast, the HSS and SIMS are characterized by a prominent disc component, contributing up to $39-98\%$ with marginal Comptonization effects ($cov_{\rm frac}\sim 0.5-38\%$) across all sources. We observe the present findings to be broadly consistent with the previously reported results for the respective sources. For example, Svoboda et al. 2024b found a steeper $\Gamma_{\rm th}\sim 4.9$ and low $cov_{\rm frac}\sim 20\%$ in the HSS of Swift J1727.816131727.816131727.8-16131727.8 - 1613, which agrees well with our findings of $\Gamma_{\rm th}\sim 4.6-5.1$ and $cov_{\rm frac}\sim 17\%$. In addition, the LHS results for Cyg X--1 (E1) are consistent with those of Krawczynski et al. (2022), except for their measurement of a high electron temperature, which is likely because of the differences in spectral energy coverage and/or the choice of model components.
@@
We also notice that the Comptonization model with a disc-like coronal geometry (comptt) yields a comparable fit to the observed spectra similar to the spherical corona model (thcomp). This evidently indicates the degeneracy in determining the corona geometry (see §4.2). Furthermore, strong reflection signatures detected in the soft state observations (HSS and SIMS) of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613, and GX $339-4$ allow the detailed investigation of alternative coronal geometries using various reflection models (see §4.2). We also observe that the reflected emission, either from a lamppost corona, modeled with relxilllp, or from a spherical corona as described in relxillCp, can be reproduced satisfactorily considering a highly ionized accretion disc ($\log\xi\sim 4$). All these findings highlight the inherent degeneracy among different disc-corona geometry models in explaining the observed spectral characteristics.

### v2

@@ 4.4 Energy-dependent Polarization
Figure 5: Variation of polarization degree (PD) with energy. In each panel, source name and spectral state (observed during different epoch) are marked. The significance of energy variation of PD for the respective cases are mentioned in units of $\sigma$ in the plot legends. The histograms represent the minimum detectable polarization at $99\%$ confidence ($\rm MDP_{99}\%$). See the text for details.
@@
We further explore energy-dependent polarization properties of the selected sources that confirm the presence of significant polarized emission in 282-8 keV energy range of IXPE. For each source except 4U $1957+115$, IGR J17091362417091-3624 and GX 3394339-4, we divide 282-8 keV energy range into six linearly spaced energy bins and estimate polarization using the model-independent approach (see also Svoboda et al. 2024a; Veledina et al. 2023). For 4U $1957+115$, IGR J17091362417091-3624 and GX 3394339-4, only four energy bins are considered for the energy-dependent study to ensure relatively improved photon statistics per energy bin (see also Marra et al. 2024). Further, we estimate the significance of the energy variation of PD over these bins following Krawczynski et al. (2022); Majumder et al. (2024b). In doing so, we consider PD in 282-8 keV energy range (see Table 5) as the null hypothesis against which the significance of energy variations in PD is computed for each epoch of a given source. With this, following Majumder et al. (2024b), we calculate the probability of rejecting the null hypothesis using (N1)(N-1) degrees of freedom, taking into account the variation of PD over $N$ energy bins for the respective cases. This essentially indicates the confidence level at which the PD variation is significant over different energy bins. The obtained significance levels of the energy variation of PD are mentioned in Table 5 for the respective cases. We observe that this significance varies from $\lesssim 1\sigma$ to $4\sigma$ across different epochs of the sources in various spectral states. In Fig. 5, we present the variation of PD with energy over different observational epochs using filled circles in the respective panels. The histograms in colors represent the $\rm MDP_{99}\%$ associated with the measurements in the selected energy bins for the corresponding epochs. Note that, for a given source, when multi-epoch observations are available, we present the results for two epochs of different spectral states only, except LMC X-3.
@@
Interestingly, we observe that Cyg X1-1 and Swift J1727.81613-1613 show similar energy variation of PD within $\sim 2-6\%$ in different spectral states (LHS, HSS, HIMS and SIMS). Moreover, for Cyg X1-1, the variation in PD with energy remains most significant at a confidence level exceeding $4\sigma$ in HSS, whereas it remains $2.8\sigma$ for Swift J1727.816131727.8-1613 in HIMS, when all observational epochs are considered (see Fig. 5). Note that, PD $\rm MDP_{99}\%$ at higher energy bins in a few epochs of several sources, indicates a null-detection. Interestingly, unlike Cyg X1-1 and Swift J1727.816131727.8-1613, a weak energy dependency of PD is noticed for IGR 17091362417091-3624 in the LHS, for which PD corresponding to two out of four energy bins remain below the $\rm MDP_{99}\%$ (see Fig. 5). We find a significant increase in PD with energy up to $\sim 11\%$ in HSS ($4\sigma$) and SPL ($4\sigma$) state of 4U 163047-47 (see Fig. 5). Similarly, distinct variations of PD over several energy bins are observed for Cyg X3-3 during LHS ($4\sigma$) and SIMS ($4\sigma$). LMC X3-3 shows constant polarization over the initial three energy bins with a sudden jump in PD up to $\sim 7.8\%$ at $\sim 5.4$ keV in both epochs (see Fig. 5). Further, we only observe a marginal variation ($1.7\sigma$) of PD in 4U 1957+115+115 up to $\sim 3\%$. Interestingly, significant ($3.3\sigma$) energy variation of polarization degree is observed for GX 3394339-4. Although, PD $\rm MDP_{99}\%$ is obtained in all four energy bins (see Fig. 5). Overall, we state that a significant ($3\sigma$) variation of polarization degree with energy is observed only for Cyg X1-1, 4U 163047-47 and Cyg X3-3, whereas a moderate ($2\sigma$ to $3\sigma$) variation is seen for Swift J1727.81613-1613 and LMC X3-3. It is worth mentioning that the energy variation of PA for all sources under consideration remains statistically insignificant ($2\sigma$). However, we note that PA drops from $30^{\circ}$ to $60^{\circ}$ at $\sim 6$ keV in 4U $1957+115$, which is marginally significant at $1.6\sigma$ confidence level.
@@ 4.5 Spectro-polarimetric Modeling
[…]
Next, we investigate the energy-dependent polarization properties using model-dependent approach for all the sources that demonstrate the signature of X-ray polarization. In doing so, we replace the polconst component with the energy-dependent polarization model polpow. Empirically, polpow assumes a power-law variation in the polarization parameters of the form PD($E$) $A_{\rm norm}\times E^{-A_{\rm index}}$ (in fraction) and PA($E$) $\psi_{\rm norm}\times E^{-\psi_{\rm index}}$ (in ) while modeling the Q and U Stokes spectra. We find that the best-fitted $\psi_{\rm index}$ remains consistent with zero within its $1\sigma$ errors and hence, we freeze it to zero during the spectral fitting of all observations. Accordingly, we obtain the best spectral fit for all sources with $\chi_{\rm red}^{2}\sim 0.97-1.46$. Furthermore, integrating PD($E$) in 282-8 keV energy range using the best-fitted model parameters ($A_{\rm index}$ and $A_{\rm norm}$), we obtain PD as $3.23-14.33\%$. Moreover, with $\psi_{\rm index}$ frozen to zero, the polarization angles are obtained as the best-fitted $\psi_{\rm norm}$ values that vary from $88.92^{\circ}$ to $85.11^{\circ}$. Note that, in a few observations of Cyg X1-1, 4U 1957+115+115 and GX 3394339-4 (see Table 6), the data is unable to constrain the parameters of polpow component even at $1\sigma$ level, and hence, we refrain from the modeling with polpow for these observations. In Table 6, we present the best-fitted model parameters obtained from the spectro-polarimetric study of all sources. We find that the model-dependent polarimetric results are in agreement with the findings of the model-independent study except Cyg X3-3, confirming the robustness of the polarization measurements for the individual sources. However, for Cyg X-3, we obtain significantly low PD compared to the results of the model-independent study, likely due to multiple low-energy unpolarized emission lines including strong iron complex around $\sim 6.4$ keV Veledina et al. (2024a), dominating Stokes spectra.
@@ 4.6 Spectro-polarimetric Correlation
Figure 6: Variation of PD (282-8 keV), obtained from model-independent analysis, as a function of the disc-to-Comptonized flux ratio ($F_{\rm ratio}$) derived in 11001-100 keV energy range from the wide-band spectral modeling. The color bar represents the covering fraction ($cov_{\rm frac}$) obtained from thcomp model, while each data point is labeled with the corresponding spectral state and observational epoch. Data points within the shaded region (light blue) are used to compute the Pearson correlation coefficient. Downward arrows indicate null-detections of polarization for the respective sources.
We examine the spectro-polarimetric correlation properties for each source based on the results obtained from wide-band spectral modeling using model M1 (see §4.2) and the polarimetric measurements (see §4.3). Specifically, we explore the correlation between the observed PD obtained from model-independent study, and best-fitted spectral parameters, namely the ratio of disc to Comptonized spectral flux ($F_{\rm ratio}$) in 11001-100 keV energy range, and the covering fraction ($cov_{\rm frac}$) obtained from the thcomp model. In Fig. 6, we present the variation of PD with $F_{\rm ratio}$ (see Table 3) using different markers, where the covering fraction ($cov_{\rm frac}$, in percent) is represented by the colors of respective markers. The color bar on the right indicates the range of $cov_{\rm frac}$. The spectral state and the corresponding epoch (in parentheses) for each observation are marked in the figure.
@@
Furthermore, we find a strong correlation (anti-correlation) between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$) with $\rho\sim 0.9$ (0.9-0.9) for Swift J1727.816131727.8-1613. In particular, a sharp drop in PD is noticed from $\sim 4.7\%$ (HIMS) to $\sim 0.4\%$ (HSS) (see also Svoboda et al. 2024b) during which $F_{\rm ratio}$ increases from $\sim 0.7$ to $\sim 66$ and $cov_{\rm frac}$ noticeably reduces down to $13\%$ from a maximum value of $82\%$. A similar trend is observed for Cyg X1-1 with $\rho=0.9$ (0.9-0.9) for the respective positive (negative) correlations between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$). As before, a low (high) $F_{\rm ratio}\sim 0.3$ ($11.5$) and high (low) $cov_{\rm frac}\sim 61-77\%$ ($0.5-31\%$) are observed in the LHS (HSS) of Cyg X1-1 with PD $\sim 3.7-4.8\%$ ($1.4-2.1\%$) (see Fig. 6). Moreover, we also find a similar behavior in the LHS of IGR J17091362417091-3624, for which a very high PD ($\sim 9\%$) is seen with $cov_{\rm frac}\sim 87\%$ and $F_{\rm ratio}\sim 0.2$. These findings suggest that a significant contribution in the polarization degree comes from the Comptonized emission for Swift J1727.816131727.8-1613, Cyg X1-1 and IGR J17091362417091-3624 in the HIMS and/or LHS. Furthermore, we notice a low polarization degree (PD $\lesssim 2\%$) for all the sources except 4U 163047-47 and LMC X3-3 in the HSS/SIMS with $F_{\rm ratio}\gtrsim 2$ and $cov_{\rm frac}\lesssim 30\%$. Interestingly, 4U 163047-47 being the source of extreme soft nature ($F_{\rm ratio}\sim 52$, $cov_{\rm frac}\sim 5\%$) manifests a high polarization degree of $\sim 8.3\%$ in the HSS, whereas it reduces to $6.8\%$ in the SPL state with $F_{\rm ratio}\sim 3.6$ and $cov_{\rm frac}\sim 38\%$. We observe that 4U $1957+115$ (GX 3394339-4) shows PD $\sim 1.2\%$ ($1.9\%$) with $cov_{\rm frac}\sim 2.7\%$ ($34\%$) and $F_{\rm ratio}\sim 18$ ($2.4$) in the HSS (SIMS). Moreover, LMC X3-3 exhibits moderate polarization with PD in the range of $2.4-3\%$ for $F_{\rm ratio}\gtrsim 4$ and $cov_{\rm frac}\sim 0.5\%$ during the HSS.
@@ 5 Discussion
In this study, we present a comprehensive spectro-polarimetric investigation of eleven BH-XRBs using quasi-simultaneous observations from IXPE, NICER, AstroSat and NuSTAR. IXPE data are analyzed to examine the polarization properties in 282-8 keV energy range, while combined NICER, AstroSat and NuSTAR observations provide insights into the spectro-temporal behavior across the wide energy range (0.51000.5-100 keV). Subsequently, the spectro-polarimetric findings are used to investigate the accretion-ejection dynamics and geometry surrounding the sources.
@@ 5.1 Spectro-temporal Properties and Spectral States
Spectro-temporal studies using NICER, NuSTAR and AstroSat observations in 0.51000.5-100 keV energy range impart distinct variability properties in various spectral states of the sources (see §4.1 and §4.2). The obtained results are in line with the spectro-temporal characteristics generally seen in the respective spectral states of BH-XRBs (Belloni et al., 2005; McClintock et al., 2009; Nandi et al., 2012; Sreehari et al., 2019b; Méndez et al., 2022; Aneesha et al., 2024). These findings reinforce the presence of distinct spectral states as observed during the multi-mission campaigns of the sources.
@@
In particular, we observe a high fractional variability amplitude $rms_{\rm tot}\%\sim 30-52$ in the PDS of LHS, whereas significantly low variability with $rms_{\rm tot}\%\sim 6-14$ are seen in the intermediate states (HIMS and SIMS) of the sources (see Table 3). Negligible power distribution at lower frequencies results in marginal variability ($rms_{\rm tot}\%\sim 1-10$) for the sources in the HSS (see Table 3 and Fig. 2). In addition, we find the presence of a strong type-C QPO feature at $\sim 1.4$ Hz (E1), which evolves to $\sim 6.7$ Hz (E2-E5), during the HIMS and SIMS of Swift J1727.81613-1613 over a wide energy range of 0.51000.5-100 keV with NICER, NuSTAR and AstroSat observations (see Fig. 2). The origin of this QPO is attributed to a strong Comptonizing corona surrounding the central source (Chakrabarti & Titarchuk, 1995; Chakrabarti et al., 2008; Nandi et al., 2024); however, based on the correlation between the evolution of QPO frequency and quasi-simultaneous radio jet ejections, Liao et al. (2024) suggest a jet-like, vertically elongated corona for Swift J1727.816131727.8-1613. A similar type-C QPO at $\sim 0.2$ Hz is detected in LHS of IGR J17091362417091-3624 (see Fig. 2). Moreover, we also observe a type-B QPO in the disc-dominated SIMS of GX 3394339-4 in the 0.5800.5-80 keV energy range from NICER and AstroSat data. Similar type-B QPOs have been observed earlier in GX 3394339-4, indicating the presence of a compact corona located very close to the black hole, accompanied by dominant disc emission (Nandi et al., 2012; Aneesha et al., 2024).
@@
The wide-band (0.7600.7-60 keV) energy spectral analysis reveals distinct spectral states of each source. The observed spectra of all sources are well described by a standard accretion disc component (diskbb) combined with thermal Comptonization of soft photons by a spherical corona (thcomp) (see Table 3 and Fig. 2). The spectral properties exhibit a strong correlation with the spectral states of the sources. In particular, during LHS and HIMS, the spectra are dominated by Comptonized emission, contributing $\sim 64-87\%$ of the total flux, accompanied by a high covering fraction ($cov_{\rm frac}\sim 61-87$). In contrast, the HSS and SIMS are characterized by a prominent disc component, contributing up to $39-98\%$ with marginal Comptonization effects ($cov_{\rm frac}\sim 0.5-38\%$) across all sources. We observe the present findings to be broadly consistent with the previously reported results for the respective sources. For example, Svoboda et al. 2024b found a steeper $\Gamma_{\rm th}\sim 4.9$ and low $cov_{\rm frac}\sim 20\%$ in the HSS of Swift J1727.816131727.8-1613, which agrees well with our findings of $\Gamma_{\rm th}\sim 4.6-5.1$ and $cov_{\rm frac}\sim 17\%$. In addition, the LHS results for Cyg X-1 (E1) are consistent with those of Krawczynski et al. (2022), except for their measurement of a high electron temperature, which is likely because of the differences in spectral energy coverage and/or the choice of model components.
@@
We also notice that the Comptonization model with a disc-like coronal geometry (comptt) yields a comparable fit to the observed spectra similar to the spherical corona model (thcomp). This evidently indicates the degeneracy in determining the corona geometry (see §4.2). Furthermore, strong reflection signatures detected in the soft state observations (HSS and SIMS) of Cyg X1-1, Swift J1727.816131727.8-1613, and GX 3394339-4 allow the detailed investigation of alternative coronal geometries using various reflection models (see §4.2). We also observe that the reflected emission, either from a lamppost corona, modeled with relxilllp, or from a spherical corona as described in relxillCp, can be reproduced satisfactorily considering a highly ionized accretion disc ($\log\xi\sim 4$). All these findings highlight the inherent degeneracy among different disc-corona geometry models in explaining the observed spectral characteristics.